The sum of the digits of a two-digit number is 9. The number is 6 times the units digit. Find the number. So far I have: a+b=9, then a+6B=n and that's as far as I can get.
Answer by Nate(3500)

The number: 10a + b
a + b = 9 which can be: a = 9 - b
and
10a + b = 6b
Plug:
10a + b = 6b
10(9 - b) = 5b
90 - 10b = 15b
90 = 15b
6 = b
Plug:
a = 9 - b = 9 - 6 = 3
Number: 10a + b = 10(3) + 6 = 30 + 6 = 36